OpenStudy (anonymous):

simplify and show work 5 1/3 + (-3 9/18)

OpenStudy (jjamz87):

retype the equation. I don't know what you're asking.

OpenStudy (ali2x2):

5 1/3 = 16/3 - 3 9/18 = -3 1/2 = - 7/2 16/3 - 7/2 = (2*16 - 7*3)/6 =(32-21)/6 = 11/6

OpenStudy (anonymous):

i agree

OpenStudy (welshfella):